Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (GMaccaferri at Wed Feb 25 09:19:29 MET 1998) Success

Torun (K. Borkowski at Wed Feb 25 09:28:10 MET 1998) Success
No known problems

Westerbork (tony foley at Wed Feb 25 12:55:24 MET 1998) Success
No known problems

Noto (C.Trigilio at Wed Feb 25 14:37:19 MET 1998) Success

Medicina (A. Polatidis at Thu Feb 26 12:56:30 MET 1998) Success
The inchword motor of HEAD 2 (the READ head) is not working. HEAD 1 (WRITE head) is positioned correctly so the data should not be affected. There are no readback tests.

Jodrell (Alastair Gunn at Thu Feb 26 15:02:45 MET 1998) Success
Strong RFI - receiver often saturated. No other known problems.

Calibration ( at Mon Mar 2 18:53:08 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name v043c.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 16:16:08 MET 1998) Minor failures
Calibration a little uncertain. Check on aspen for updated values in v043ccal.eb.Z